A car moving with velocity of 36km/h accelerates uniformly at 2m/s² until it reaches a velocity of 54 km/h.the time taken is ?
1
Expert's answer
2020-12-06T17:20:41-0500

The velocity "v" of an object moving with a constant acceleration at any moment of time "t" is given by the following formula:



"v = v_0+at"


where "v_0 =36km\/h = 10m\/s" is the initial velocity of the car and "a = 2m\/s^2" is its acceleration. If we set "v = 54km\/h = 15m\/s", we can find the time that was taken to come from "v_0" to "v":



"t = \dfrac{v-v_0}{a} = \dfrac{15-10}{2} = 2.5s"

Answer. 2.5 s.